OpenStudy (anonymous):

Will medal, Factor the trinomial 36x^2+24x+4 A: (36x+4)^2 B: (18x+2)^2 C: (6x+2^2 D: (6x+4)^2 E: none

OpenStudy (cwrw238):

take the square root of the first and last term

OpenStudy (anonymous):

36x and 2?

OpenStudy (jhannybean):

Factor out a 4. \[4(9x^2+6x+1)\]\[= 4(x^2+6x+9)\]\[= 4(x+3)(x+3)\]\[= 4\left(x+\frac{3}{9}\right)\left(x+\frac{3}{9}\right)\]\[= 4\left(x+\frac{1}{3}\right)\left(x+\frac{1}{3}\right)\]\[= 4(3x+1)(3x+1)\]

OpenStudy (jhannybean):

So what I did was: 1. Factor out the LCM 4 2. factored what was inside the parenthesis. 3. Found 2 numbers that multiplied to give me 9, added to give me 6. 4. Because I had originally multiplied the 9 over to the constant to give me a simpler quadratic equation to solve for, I had to readd it back into my equation, so I divided all my terms by it. 5. simplified the fraction. 6. Multiplied the denominator over to the variable to simplify the function over a common denominator.

OpenStudy (anonymous):

okay i understand that

OpenStudy (jhannybean):

Oh you are the first!`

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

jhanny its the ancient arabic method that idiotic modern men have found on a tablet in a cave, and claimed it as their own

OpenStudy (jhannybean):

Yes, what most people do not understand is that when you take a complex quadratic function, such as \(9x^2+6x+1\) you can simplify it by multiplying the leading coefficient, \(9\) to the constant at the end, \(1\). This allows for easier factoring. because you're trying to keep to the ORIGINAL equation, you must divide by the same leading coefficient \(9\) when attaining the factored form.

OpenStudy (anonymous):

Yeah i didnt know that untill you explained it but i understood

OpenStudy (anonymous):

in sorry what letter was that?

OpenStudy (jhannybean):

What letter?

OpenStudy (anonymous):

i had a-e, and sorry for typeing funny i cup open my hand

OpenStudy (anonymous):

cut*

OpenStudy (jhannybean):

Im not seeing the answer choice there.

OpenStudy (anonymous):

A: (36x+4)^2 B: (18x+2)^2 C: (6x+2^2 D: (6x+4)^2 E: none

OpenStudy (perl):

@Jhannybean are all these steps correct/intentional which you wrote 4( 9x^2 + 6x + 1) = 4( x^2 + 6x + 9) = 4(x+3)(x+3) = 4( x + 3/9) (x+3/9) =4( x + 1/3) (x+1/3) = 4 ( 3x+1)(3x+1)

OpenStudy (perl):

this method is very clever :) I agree with Jhanny, i dont see answer choice

OpenStudy (anonymous):

thank you both can you help with another?

OpenStudy (perl):

ok

OpenStudy (jhannybean):

Nevermind, I am wrong, it is there.

OpenStudy (anonymous):

o.o?

OpenStudy (anonymous):

how so?

OpenStudy (perl):

4 ( 3x+1)^2 = 2 (3x+1) * 2(3x+1)

OpenStudy (perl):

4 ( 3x+1)^2 = (6x + 2) * (6x+2)

OpenStudy (perl):

but the factorization was correct

OpenStudy (jhannybean):

I did not know how to explain it.

OpenStudy (anonymous):

Thats okay

OpenStudy (perl):

4( 9x^2 + 6x + 1) = 4( x^2 + 6x + 9) = 4(x+3)(x+3) = 4( x + 3/9)(x+3/9) =4(x + 1/3)(x+1/3) = 4(3x+1)(3x+1) =2*2 (3x+1)*(3x+1) = 2(3x+1) * 2(3x+1) = (6x + 2)( 6x+2) =(6x+2)^2

OpenStudy (jhannybean):

Oh now that makes sense. \[4(3x+1)^2\]\[2^2(3x+1)^2\]\[\left[2(3x+1)\right]^2\]\[2(3x+1)\cdot 2(3x+1)\]

OpenStudy (jhannybean):

\[=(6x+2)(6x+2)\]\[=(6x+2)^2\]

OpenStudy (anonymous):

Thank you both so much this really helps, could i get help with this one x^2-17x+16 A:(x-2)(x-8) B:(x-16)(x-1) C:(x-16)(x+1) D:(x-4)(x+4) E: none

OpenStudy (jhannybean):

What are two numbers that multiply to give you 16, add to give you -17?

OpenStudy (anonymous):

umm 8*2 is 16 4*4 is 16

OpenStudy (jhannybean):

Think bigger, and take the factors of 17 into consideration.

OpenStudy (cwrw238):

answer C would be Ok for the first one - it looks like an end bracket is missing

OpenStudy (cwrw238):

it should be (6x + 2)^2

OpenStudy (anonymous):

Im really sorry i have to go get lunch and dinner for my family thank you for all the help @Jhannybean

OpenStudy (jhannybean):

No problem :)

OpenStudy (cwrw238):

Hi jhannybean - you have made 2 fans here!

OpenStudy (cwrw238):

I tried to message you back but could not

OpenStudy (jhannybean):

Oh, because I have been getting spammed quite a bit lately.